OpenStudy (anonymous):

At the start of a basketball game, the referee tosses a ball into the air for the jump ball. The equation that models the height of the ball is h(t) = -16t2 + 30t + 8. During what time interval is the height of the ball above 10 feet? (Round your answer to the nearest tenth of a second.)

OpenStudy (anonymous):

they need to vary these questions solve \[-16t^2+30t+8>10\] \[-16t^2+30t-2>0\] \[8t^2-15t+2<0\]

OpenStudy (anonymous):

find the zeros. your answer will be the interval between them

OpenStudy (anonymous):

zeros are \[\frac{15-\sqrt{161}}{16}\] and its conjugate \[\frac{15+\sqrt{161}}{16}\]

OpenStudy (anonymous):

use a calculator to find the decimal approximation if that is what you need

OpenStudy (anonymous):

from .1444 to 1.73 rounded

OpenStudy (anonymous):

thank you